OpenStudy (anonymous):

2cos^2x-1=0 find all solutions (there can be more than 1) pi/8 -7pi/4 15pi/4 3pi/4

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \bf 2cos^2x-1=0 }\) \(\Large\color{blue}{ \bf 2cos^2x=1 }\) \(\Large\color{blue}{ \bf cos^2x=1/2 }\) \(\Large\color{blue}{ \bf cosx=1/4 }\)

OpenStudy (aric200):

2 * cos² x - 1 = 0 <=> 2 * cos² x = 1 <=> cos² x = 1/2 <=> cos x = - sqr (1/2) or cos x = sqr (1/2) <=> cos x = - sqr (2) / 2 or cos x = sqr (2) / 2 <=> x = 3pi/4 or x = 5pi/4 or x = 7pi/4 or x = pi/4 Answers : pi/4 ; 3pi/4 ; 5pi/4 ; 7pi/4 ===> radians 45 ; 135 ; 225 ; 315 ===> degrees *******